OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (2, 2) (3, 3)

OpenStudy (anonymous):

@AfifaRahman

OpenStudy (anonymous):

m= (y1-y2)/(x1-x2)

OpenStudy (anonymous):

m= (2-3)/(2-3)=1

OpenStudy (anonymous):

Now the equation is y=x+b

OpenStudy (anonymous):

but in the coordinates y=x which means x=x+b which gives b=0

OpenStudy (anonymous):

hence the equation is y=x

OpenStudy (afifarahman):

sorry anna, my internet connection was down, its answer is \[y=x\]